### v1

@@ 4.2. Does the RL manager implicitly predict recommendations of team members?
If the RL manager learned chess, did it simply learn to become an expert manager? Recall that unlike the expert manager, the RL manager was never shown the move recommendations of Maia or Leela. To become like the expert manager, it would have to implicitly learn to predict the move recommendations.
@@
Figure 7. Top - attention to origin squares, Bottom - attention to destination squares. No major difference is observed compared to the controls. Whiskers in the box plots represent farthest data point lying within 1.5x the inter-quartile range from the box.
We thus find no evidence that the network learns to recognize Maia or Leela moves. This indicates that relative advantages can be ascertained with a more abstract understanding of the qualities of the team members. It would also imply that whatever domain knowledge was accumulated by the RL manager, was not being used to judge the move recommendations on their merits, as the expert manager does.
@@ 4.3. Are learned relative advantages interpretable?
It is interesting to test whether the RL manager has learned to distinguish relative advantages between Maia and Leela, using human-understandable features, such as those found in the past between humans and older chess programs by (Kopec and Bratko,, 1982). This would be informative regarding humans and machines, and it might allow more efficient production of synergy using a heuristic manager instead. To explore this possibility, we formulated a set of human understandable features of the chess board, based on strategic chess concepts, and extracted them for positions in which Maia and Leela were assigned advantage by the manager. These include:
@@
Differences in mean feature values are summarized in table 1. We again measure effect size using non-parametric $A_{w}$ (Li et al.,, 2021). Effect size is always shown in the direction of lower to higher mean value. A value of 0.5 represents no effect, 1 represents highest effect. We find that effect sizes are all small. It does not seem that the RL manager uses human understandable features.
@@ 5. Addendum: Maia Human Likeness
While not essential for our conclusions, we assumed in this work that Maia is a reasonable proxy for humans. It has over 50% accuracy in matching human moves (chess has over 30 available moves per turn on average), more than other chess engines (McIlroy-Young et al.,, 2020). However, we deploy Maia in a team, potentially leading it to out of distribution positions. To test if this degrades its human likeness, we check for human biases reflected in Maia’s playing style, when deployed in a team. (Afek and Neiman,, 2012) documented ”invisible moves”, i.e. positions in which humans miss good moves due to bias, such as:
@@
  • ”Didactic bias” - a bias to follow heuristic rules taught in chess clubs. These include trying to control the center, castling early and developing major pieces.
@@
Figure 8. Move feature frequencies compared between Maia (blue) and Leela (orange) recommendations. Error bars represent SEM.
@@ 6. Related work
Teams have been known to achieve synergy at various tasks, i.e. better performance than any member of the team. This phenomenon has been studied in both human teams and machine ensembles and there is growing work on human-machine teams (Wilder et al.,, 2020; Rosenfeld et al.,, 2017; Kim et al.,, 2017). In both human-team and machine-ensemble studies, diversity in team composition has been shown to correlate with higher team performance, also known as the ”diversity bonus” (Page,, 2019). This is encouraging for human-machine teams which certainly have cognitive diversity.
@@
However, not all teams achieve synergy. In human teams, factors for success include effective communication between team members (Malone,, 2019). This may be complicated in human-machine teams due to lack of explainability of machine models. Collective decision-making has also been studied analytically. The classic Jury theorem (de Condorcet,, 1785) posits that majority vote leads to correct decisions more often than for any of the voters individually, as long as each voter is correct at least 50% of the time. However, majority vote is not always applicable, for example if there are only two agents, as in our case. And it fails when some voters are correct less than 50% of the time, making voting innapropriate for asymmetric teams.
@@
In some cases it is possible to judge the recommendations of team members based on their merit. For example, (Osborne and Rubinstein,, 1997) studied boundedly-rational agents as participants in ”machine games”. In machine games, players select from among a set of imperfect ”machines” to play on their behalf in a repeated base-game. Our work can be viewed as a type of machine game where one base player is human-like and one is an RL trained network.
@@
But choosing based on merit is problematic when there is difficulty verifying the utility of the recommendations. (Amir et al.,, 2018) for example found that human teams failed at solving some types of analytical problems, even when individual team members succeeded. This occurred when the correct solution was not easily verifiable and the other team members voted it down. We study collective intelligence in a low-verifiability regime (long horizon sequential decision-making). As we saw, finding a synergetic combination of Maia and Leela moves required a non-trivial process.
@@
Mixture of Experts (Yuksel et al.,, 2012) is a successful learning paradigm, that trains a ”gating mechanism” or ”router” (which we denoted ”manager”) to select from a set of parallel learners. Our work capitalizes on these concepts to model the human-machine team. But our setting is bottom-up, taking pretrained frozen policies, and then training only the ”manager”.
@@
To the best of our knowledge, MoE has not been directly applied to model human-machine teaming. Other paradigms have been explored. For example, (Wilder et al.,, 2020) advocated end-to-end learning for machine models together with humans. Our work by contrast, considers how to leverage pretrained machine models in teams, even though they were not specifically developed for a team setting. (Kim et al.,, 2017) envisioned a division of labor, whereby high-level planning is done by humans, which is then fed to machines for subsequent low-level optimization by the machine. By contrast, our work considers when human and machine team members recommend actions in parallel. (Irving et al.,, 1018) proposes a debating mechanism between machines with a human arbitrator for safe AI. Our work examines how far one can go without information exchange or deliberation between agents. (Rosenfeld et al.,, 2017) developed an agent to assist a human manager in operating multiple robots. Our work shares similar motivations.
@@
Chess has been referred to as the ”drosophila of AI” (McCarthy,, 1990). Its simple setup on the one hand, and complex state-space on the other hand, makes it attractive also for studying the human-machine teaming problem.
@@
In chess, (Krakowski et al.,, 2022) analyzed the relative performance of a set of players in human tournaments, and the same players as centaurs in freestyle tournaments. They found no correspondence between the two, i.e. humans don’t have to be top players in chess to be effective centaurs. They conclude from this that in order to succeed as centaur players, an entirely different set of skills is required than domain knowledge in chess. They speculated that these might be technical skills such as data analysis or configuring the machine.
@@
We show that there exists significant potential for synergy in centaur teams by contributing substantive human or humanlike judgement, as with the Maia engine, even if the partner machine is vastly superior. However, we also show that learning these relative advantages becomes difficult as asymmetry of the team increases. Furthermore, we find indications that the function of recognizing relative advantages between team members requires some domain knowledge, even if not superior domain knowledge. This might indicate that a human centaur player with no knowledge of chess would probably fail to achieve synergy.
@@
Attempts to characterize human-machine relative advantages in chess is not new. Last century, Bratko and Kopec (Kopec and Bratko,, 1982) devised a set of tests on different types of positions, and found that while machines excel at solving ”tactical” problems, humans are better at ”positional” problems, particularly regarding pawn structures. We conducted a Bratko-Kopec-like experiment for the deep-learning era and found that traditional strategic concepts in chess no longer explain relative advantages (Section 4.3).
@@
More recently, (Wang et al.,, 2023) trained an adversarial network to beat a superhuman network in the game go, but then showed that the adversarial policy was trivially defeatable by humans. That is, they produced a violation of transitivity in game skill between a superhuman policy, an adversarial policy and humans. Their work illustrates that even superhuman networks can develop blind spots. In fact, there are collections of peculiar chess problems that are difficult for superhuman chess programs, even though humans can reason out their solutions (Steingrimsson,, 2021). (Zahavy et al.,, 2023) considered a method to improve machine performance on these problems by constructing a team of diverse policies. Our work leverages human/machine diversity specifically.
@@
Behavioral clones are a popular method for generating scaled synthetic data on human behavior to train collaborative agents (Kantack, N.,, 2021). Here we use them to test for synergy in a human-machine team.
@@ 7.1. Contributions
Firstly, we demonstrate (with the oracle manager) that there is potential for substantial human-machine synergy, which persists even in highly asymmetric teams. However, in practice finding the relative advantages that lead to this synergy is non-trivial. We propose modeling the human-machine team as a Mixture of Experts to isolate the function of finding relative advantages. Practically, a MoE model for human-machine teams could inform team decisions with a ”relative advantage score”. This might be better than using, for example, confidence scores, for which neural networks are not well callibrated (Guo et al.,, 2017), and which are not comparable between different agents.
We found that a subject matter expert can find a moderate portion of the potential relative advantages and that increasing expertise has quickly diminishing returns for this task. The limitations of subject matter expertise are probably due to the ”curse of knowledge”, i.e. the tendency of the expert to overestimate the team members when planning ahead. This might have implications for debates about the role of subject matter expertise in management.
We also found that relative advantages between agents can be learned, without communication or explainability of their recommendations, in both symmetric and moderately asymmetric teams. This is encouraging for humans teaming with strong but unexplainable machine models. Furthermore, we find (by examining the network of the RL manager) that identifying relative advantages requires at least some minimal domain knowledge, even if not superior domain knowledge to the team members. This means that human operators should not forego obtaining domain knowledge, even when machines reach higher levels of proficiency in a task.
@@
To reach our conclusions, we made use of a behavioral clone - the Maia network (McIlroy-Young et al.,, 2020). Our methodology expands the use of behavioral clones, from serving to train collaborative agents, to testing human-machine synergy.

### v2

@@ 4.2. Does the RL manager implicitly predict recommendations of team members?
If the RL manager learned some chess, did it inadvertently become an expert manager? For this to happen, it would have to learn to predict the move recommendations, since it is not given them.
@@
Figure 7. Comparison of attention to moves. Top - attention to origin squares, Bottom - attention to destination squares. No major difference is observed compared to the controls. Whiskers in the box plots represent farthest data point lying within 1.5x the inter-quartile range from the box.
[…]
Investigation of attention scores to moves by different players
We thus find no evidence that the network learns to recognize Maia or Leela moves. This implies that whatever domain knowledge was accumulated by the RL manager, it was not being used to judge the move recommendations on their merits, as the expert manager does. Instead, relative advantages are ascertained with a more abstract understanding of the qualities of the team members.
@@ 4.3. Are learned relative advantages interpretable?
It is interesting to test whether the RL manager has learned to distinguish relative advantages between Maia and Leela using human-understandable features, such as those found in the past between humans and older chess programs by Kopec and Bratko (1982). This would be informative regarding relative advantages of humans and machines, and it might allow more efficient production of synergy using a rule-based manager instead. To explore this possibility, we formulated a set of human understandable features of the chess board, based on strategic chess concepts. These include:
@@
Taking a set of board positions in which the RL manager gave preference to the Maia or Leela move, we extracted the above features. We then compared the mean value for each feature in positions where Maia was given preference, to the mean value when Leela was given preference. Results are summarized in table 1. We again measure effect size using non-parametric $A_{w}$ Li et al. (2021). Effect size is always shown in the direction of lower to higher mean value. A value of 0.5 represents no effect, 1 represents highest effect. We find that effect sizes are all small. This implies that no one feature by itself seems to be relied upon by the RL manager to find synergy.
This doesn’t mean that the RL manager doesn’t implicitly use some combination of these features. To test for this, we distilled the transformer based manager into a fully connected deep neural network (20 layers of width 256), which takes the above features as input, and was trained to predict the output of the transformer network. Using the distilled model as manager, it managed to achieve significant synergy in the symmetric team over the test opening positions, with WDL=0.47. This is less than the transformer model, but significantly better than Maia alone (Z=4.77). This means that a combination of the features is relevant for synergy.
However, deep neural networks combine features in a complex non-linear fashion, making interpretability hard. To investigate whether a linear combination of the features might be useful for producing synergy, we used the same RL training regime outlined in section 2.3 above to train a logistic regression model. The model converged to a manager that chooses the same player every time and so does not achieve synergy. We conclude that a complex non-linear combination of the features is required to achieve synergy.
@@ 5. Addendum: Maia Human Likeness
While not essential for our conclusions, we assumed in this work that Maia is a reasonable proxy for humans. It has over 50% accuracy in matching human moves (chess has over 30 available moves per turn on average), more than other chess engines McIlroy-Young et al. (2020). However, we deploy Maia in a team, potentially leading it to out of distribution positions. To test if this degrades its human likeness, we check for human biases reflected in Maia’s playing style, when deployed in a team. Afek and Neiman (2012) documented ”invisible moves”, i.e. positions in which humans miss good moves due to bias, such as:
@@
  • ”Didactic bias” - a bias to follow heuristic rules taught in chess clubs. These include trying to control the center, castling and developing major pieces.
@@
Figure 8. Move feature frequencies compared between Maia and Leela recommendations. Error bars represent SEM.
[…]
Comparison of feautures in Maia and Leela moves
@@ 6. Related work
Teams have been known to achieve synergy at various tasks, i.e. better performance than any member of the team. This phenomenon has been studied in both human teams and machine ensembles and there is growing work on human-machine teams Wilder et al. (2020); Rosenfeld et al. (2017); Kim et al. (2017). In both human-team and machine-ensemble studies, diversity in team composition has been shown to correlate with higher team performance, also known as the ”diversity bonus” Page (2019). This is encouraging for human-machine teams which certainly have cognitive diversity.
@@
However, not all teams achieve synergy. In human teams, factors for success include effective communication between team members Malone (2019). This may be complicated in human-machine teams due to lack of explainability of machine models. Collective decision-making has also been studied analytically. The classic Jury theorem de Condorcet (1785) posits that majority vote leads to correct decisions more often than for any of the voters individually, as long as each voter is correct at least 50% of the time. However, majority vote is not always applicable, for example if there are only two agents, as in our case. And it fails when some voters are correct less than 50% of the time, making voting inappropriate for highly asymmetric teams.
@@
In some cases it is possible to judge the recommendations of team members based on their merit. For example, Osborne and Rubinstein (1997) studied boundedly-rational agents as participants in ”machine games”. In machine games, players select from among a set of imperfect ”machines” to play on their behalf in a repeated base-game. Our work can be viewed as a type of machine game where one base player is human-like and one is an RL trained network.
@@
But choosing based on merit is problematic when there is difficulty verifying the utility of the recommendations. Amir et al. (2018) for example found that human teams failed at solving some types of analytical problems, even when individual team members succeeded. This occurred when the correct solution was not easily verifiable and the other team members voted it down. We study collective intelligence in a low-verifiability regime (long horizon sequential decision-making). As we saw, finding a synergetic combination of Maia and Leela moves required a non-trivial process.
@@
Mixture of Experts Yuksel et al. (2012) is a successful learning paradigm, that trains a ”router” (which we denoted ”manager”) to select from a set of parallel learners. Our work capitalizes on these concepts to model the human-machine team. But our setting is bottom-up, taking pretrained frozen policies, and then training only the ”manager”.
@@
To the best of our knowledge, MoE has not been directly applied to model human-machine teaming. Other paradigms have been explored. For example, Wilder et al. (2020) advocated end-to-end learning for machine models together with humans. Our work by contrast, considers how to leverage pretrained machine models in teams, even though they were not specifically developed for a team setting. Kim et al. (2017) envisioned a division of labor, whereby high-level planning is done by humans, which is then fed to machines for subsequent low-level optimization by the machine. By contrast, our work considers when human and machine team members recommend actions in parallel. Irving et al. (2018) proposes a debating mechanism between machines with a human arbitrator for safe AI. Our work examines how far one can go without information exchange or deliberation between agents. Rosenfeld et al. (2017) developed an agent to assist a human manager in operating multiple robots. Our work shares similar motivations.
@@
Chess has been referred to as the ”drosophila of AI” McCarthy (1990). Its simple setup on the one hand, and complex state-space on the other hand, makes it attractive also for studying the human-machine teaming problem.
@@
In chess, Krakowski et al. (2022) analyzed the relative performance of a set of players in human tournaments, and the same players as centaurs in freestyle tournaments. They found no correspondence between the two, i.e. humans don’t have to be top players in chess to be effective centaurs. They conclude from this that in order to succeed as centaur players, an entirely different set of skills is required than domain knowledge in chess. They speculated that these might be technical skills such as data analysis or configuring the machine.
@@
We show that there exists significant potential for synergy in centaur teams by contributing substantive human or humanlike judgement, as with the Maia engine, even if the partner machine is vastly superior. However, we also show that learning these relative advantages becomes difficult as asymmetry of the team increases. Furthermore, we find indications that the function of successfully learning to recognize relative advantages between team members, requires at least some domain expertise. This might indicate that a human centaur player with no knowledge of chess would probably fail to achieve synergy.
@@
Attempts to characterize human-machine relative advantages in chess are not new. Last century, Bratko and Kopec (Kopec and Bratko, 1982) devised a set of tests on different types of positions, and found that while machines excel at solving ”tactical” problems, humans are better at ”positional” problems, particularly regarding pawn structures. We conducted a Bratko-Kopec-like experiment for the deep-learning era and found that traditional strategic concepts in chess no longer explain relative advantages (Section 4.3).
@@
More recently, Wang et al. (2023) trained an adversarial network to beat a superhuman network in the game go, but then showed that the adversarial policy was trivially defeatable by humans. That is, they produced a violation of transitivity in game skill between a superhuman policy, an adversarial policy and humans. Their work illustrates that even superhuman networks can develop blind spots. In fact, there are collections of peculiar chess problems that are difficult for superhuman chess programs, even though humans can reason out their solutions Steingrimsson (2021). Zahavy et al. (2023) considered a method to improve machine performance on these problems by constructing a team of diverse policies. Our work leverages human/machine diversity specifically.
@@
Behavioral clones are a popular method for generating scaled synthetic data on human behavior to train collaborative agents Kantack (2021). Here we use them to test for synergy in a human-machine team.
@@ 7.1. Contributions
Firstly, we demonstrate (with the oracle manager) that there is potential for substantial human-machine synergy, which persists even in highly asymmetric teams. However, in practice finding the relative advantages that lead to this synergy is non-trivial. We propose modeling the human-machine team as a Mixture of Experts to isolate the function of finding relative advantages. Practically, a MoE model for human-machine teams, could inform team decisions with a ”relative advantage score”. This might be better than using, for example, confidence scores, for which neural networks are not well calibrated (Guo et al., 2017), and which are not comparable between different kinds of agents.
We found that a subject matter expert can find a moderate portion of the potential relative advantages and that increasing expertise has quickly diminishing returns for this task. The limitations of subject matter expertise are probably due to the ”curse of knowledge”, i.e. the tendency of the expert to overestimate the team members when planning ahead.
We also found that relative advantages between agents can be learned, without communication or explainability of their recommendations, in both symmetric and moderately asymmetric teams. This is encouraging for humans teaming with strong but unexplainable machine models. Furthermore, we find (by examining the network of the RL manager) that identifying relative advantages requires at least some minimal domain knowledge, even if it is not superior to that of the team members. This means that human operators should not forego obtaining domain knowledge, even when machines reach higher levels of proficiency.
@@
To reach our conclusions, we made use of a behavioral clone - the Maia network McIlroy-Young et al. (2020). Our methodology expands the use of behavioral clones, from serving to train collaborative agents, to testing human-machine synergy.